# A Continuous-Time Ensemble Kalman–Bucy Smoother for Causal Inference and Model Discovery

### 28 Apr 2026 v1

@@ Appendix B Moment Consistency of the EnKBS in the Linear-Gaussian Regime
Rearranging (34) gives $\mathbb{E}\big[\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}-\mathbf{\hat{x}}_{\mathrm{f}}^{(i)})^{\top}\big]=\mathbf{0}$, meaning RTS smoother covariance consistency is achieved if the smoother decreases the ensemble posterior uncertainty beyond the filter solution along directions where the smoother anomaly gain, $\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}-\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}$, is uncorrelated or orthogonal to the filter anomaly $\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}$.
We now show that (34) holds when the stochastic EnKBF formulation is adopted. For the stochastic EnKBF (7), the filter anomalies evolve according to
$\mathrm{d}\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}=(\mathbf{F}-\mathbf{K}\mathbf{H})\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}-\mathbf{K}\mathbf{\Gamma}^{1/2}\,\mathrm{d}\mathbf{W}^{(i)},$ (35)
where $\mathbf{K}=\mathbf{P}_{\mathrm{f}}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}$. Recall that both the filter anomaly (35) and the smoother anomaly (31) are driven by the exact same system noise $\mathrm{d}\mathbf{B}^{(i)}$. We apply Itô’s product rule to $\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}$, where the independent observation noise $\mathrm{d}\mathbf{W}^{(i)}$ is uncorrelated with $\mathrm{d}\mathbf{B}^{(i)}$ and the shared system noise generates a non-zero Itô correction term $\mathbb{E}\big[(\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)})(\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)})^{\top}\big]=\mathbf{\Sigma}\,\mathrm{d}t$. Taking then an expectation, we obtain the evolution equation for the cross-covariance $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbb{E}\big[\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big]$:
$\dot{\mathbf{P}}_{\mathrm{f},\mathrm{s}}$ $=\mathbb{E}\left[(\mathbf{F}-\mathbf{K}\mathbf{H})\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}(\mathbf{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\right]+\mathbb{E}\left[\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}\Big(\big(\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big)\mathbf{\hat{x}}_{\mathrm{s}}^{(i)}-\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}\Big)^{\top}\right]+\mathbf{\Sigma}$
@@
On the other hand, the standard continuous-time Riccati equation for the filter covariance $\mathbf{P}_{\mathrm{f}}$ can be written as [53, 34]
@@
Comparing (36) and (37), both $\mathbf{P}_{\mathrm{f},\mathrm{s}}$ and $\mathbf{P}_{\mathrm{f}}$ satisfy the ODE:
@@
Since the smoother ensemble is “initialized” identically to the filter one at the terminal time $t=T$, we have $\mathbf{P}_{\mathrm{f},\mathrm{s}}(T)=\mathbf{P}_{\mathrm{f}}(T)$. Therefore, under appropriate regularity conditions, the solution of (38) is unique and $\mathbf{P}_{\mathrm{f},\mathrm{s}}\equiv\mathbf{P}_{\mathrm{f}}$ for all $t$.
Substituting $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbf{P}_{\mathrm{f}}$ back into (33), the smoother covariance equation simplifies to
@@
We emphasize that recovering the RTS smoother covariance equation relies on the stochastic EnKBF formulation (9). For other deterministic EnKBF formulations based on ensemble transformations [81, 82], $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbf{P}_{\mathrm{f}}$ does not generally hold. This theoretical mismatch is caused by the coupling with the filter member $\mathbf{x}_{\mathrm{f}}^{(i)}$ in the EnKBS formulation (9). For example, consider a continuous-time deterministic EnKBF formulation [60, 57]:
$\mathrm{d}\mathbf{x}^{(i)}_{\mathrm{f}}=\mathbf{F}\mathbf{x}^{(i)}_{\mathrm{f}}\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}+\mathbf{P}_{\mathrm{f}}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\left(\,\mathrm{d}\mathbf{y}-\frac{\mathbf{H}\mathbf{x}^{(i)}_{\mathrm{f}}+\mathbf{H}\overline{\mathbf{x}}_{\mathrm{f}}}{2}\,\mathrm{d}t\right).$ (40)
Its anomalies obey the dynamics with a half-gain drift (cf. (35)):
$\mathrm{d}\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}=\left(\mathbf{F}-\frac{1}{2}\mathbf{K}\mathbf{H}\right)\mathbf{\hat{x}}_{\mathrm{f}}^{(i)}\,\mathrm{d}t+\mathbf{\Sigma}^{1/2}\,\mathrm{d}\mathbf{B}^{(i)}.$ (41)
@@
Clearly, (42) does not match the continuous-time Riccati equation (37) for the filter covariance $\mathbf{P}_{\mathrm{f}}$. To see the discrepancy explicitly, we define the difference matrix $\mathbf{D}:=\mathbf{P}_{\mathrm{f},\mathrm{s}}-\mathbf{P}_{\mathrm{f}}$. Subtracting (37) from (42) yields
@@
Although $\mathbf{D}(T)=\mathbf{0}$, the positive semi-definite term $\frac{1}{2}\mathbf{P}_{\mathrm{f}}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\mathbf{H}\mathbf{P}_{\mathrm{f}}$ generally forces $\mathbf{D}(t)\neq\mathbf{0}$ for $t<T$. Hence, $\mathbf{P}_{\mathrm{f},\mathrm{s}}\neq\mathbf{P}_{\mathrm{f}}$ for this deterministic EnKBF. Consequently, the resulting smoother ensemble covariance $\mathbf{P}_{\mathrm{s}}$ will deviate from the RTS smoother covariance equation (39).
Nevertheless, despite this theoretical mismatch, we have tested this deterministic EnKBF (40) as a forward pass in our framework. It yields results similar to those of the stochastic EnKBF (7) across all numerical experiments presented in this paper (not shown). In addition, for either choice of the forward EnKBF, the smoother consistently improves upon the filter by reducing both the estimation error and the ensemble spread (with behavior similar to that shown in Figure 6).

### 03 May 2026 v2

@@ Appendix B Moment Consistency of the EnKBS in the Linear-Gaussian Regime
Rearranging Eq. (34) gives $\mathbb{E}\big[\vec{\hat{x}}_{\mathrm{f}}^{(i)}(\vec{\hat{x}}_{\mathrm{s}}^{(i)}-\vec{\hat{x}}_{\mathrm{f}}^{(i)})^{\top}\big]=\vec{0}$, meaning RTS smoother covariance consistency is achieved if the smoother decreases the ensemble posterior uncertainty beyond the filter solution along directions where the smoother anomaly gain, $\vec{\hat{x}}_{\mathrm{s}}^{(i)}-\vec{\hat{x}}_{\mathrm{f}}^{(i)}$, is uncorrelated or orthogonal to the filter anomaly $\vec{\hat{x}}_{\mathrm{f}}^{(i)}$.
We now show that Eq. (34) holds when the stochastic EnKBF formulation is adopted. For the stochastic EnKBF Eq. (7), the filter anomalies evolve according to
$\mathrm{d}\vec{\hat{x}}_{\mathrm{f}}^{(i)}=(\mathbf{F}-\mathbf{K}\mathbf{H})\vec{\hat{x}}_{\mathrm{f}}^{(i)}\mathrm{d}t+\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)}-\mathbf{K}\mathbf{\Gamma}^{1/2}\mathrm{d}\vec{W}^{(i)},$ (35)
where $\mathbf{K}=\mathbf{P}_{\mathrm{f}}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}$. Recall that both the filter anomaly Eq. (35) and the smoother anomaly Eq. (31) are driven by the exact same system noise $\mathrm{d}\vec{B}^{(i)}$. We apply Itô’s product rule to $\vec{\hat{x}}_{\mathrm{f}}^{(i)}(\vec{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}$, where the independent observation noise $\mathrm{d}\vec{W}^{(i)}$ is uncorrelated with $\mathrm{d}\vec{B}^{(i)}$ and the shared system noise generates a non-zero Itô correction term $\mathbb{E}\big[(\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)})(\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)})^{\top}\big]=\mathbf{\Sigma}\mathrm{d}t$. Taking then an expectation, we obtain the evolution equation for the cross-covariance $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbb{E}\big[\vec{\hat{x}}_{\mathrm{f}}^{(i)}(\vec{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\big]$:
$\dot{\mathbf{P}}_{\mathrm{f},\mathrm{s}}$ $=\mathbb{E}\left[(\mathbf{F}-\mathbf{K}\mathbf{H})\vec{\hat{x}}_{\mathrm{f}}^{(i)}(\vec{\hat{x}}_{\mathrm{s}}^{(i)})^{\top}\right]+\mathbb{E}\left[\vec{\hat{x}}_{\mathrm{f}}^{(i)}\Big(\big(\mathbf{F}+\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\big)\vec{\hat{x}}_{\mathrm{s}}^{(i)}-\mathbf{\Sigma}\mathbf{P}^{-1}_{\mathrm{f}}\vec{\hat{x}}_{\mathrm{f}}^{(i)}\Big)^{\top}\right]+\mathbf{\Sigma}$
@@
On the other hand, the standard continuous-time Riccati equation for the filter covariance $\mathbf{P}_{\mathrm{f}}$ can be written as (Kalman and Bucy, 1961; Jazwinski, 1970)
@@
Comparing Eq. (36) and Eq. (37), both $\mathbf{P}_{\mathrm{f},\mathrm{s}}$ and $\mathbf{P}_{\mathrm{f}}$ satisfy the ODE:
@@
Since the smoother ensemble is “initialized” identically to the filter one at the terminal time $t=T$, we have $\mathbf{P}_{\mathrm{f},\mathrm{s}}(T)=\mathbf{P}_{\mathrm{f}}(T)$. Therefore, under appropriate regularity conditions, the solution of Eq. (38) is unique and $\mathbf{P}_{\mathrm{f},\mathrm{s}}\equiv\mathbf{P}_{\mathrm{f}}$ for all $t$.
Substituting $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbf{P}_{\mathrm{f}}$ back into Eq. (33), the smoother covariance equation simplifies to
@@
We emphasize that recovering the RTS smoother covariance equation relies on the stochastic EnKBF formulation Eq. (7). For other deterministic EnKBF formulations based on ensemble transformations (Amezcua et al., 2014; de Wiljes et al., 2020), $\mathbf{P}_{\mathrm{f},\mathrm{s}}=\mathbf{P}_{\mathrm{f}}$ does not generally hold. This theoretical mismatch is caused by the coupling with the filter member $\vec{x}_{\mathrm{f}}^{(i)}$ in the EnKBS formulation Eq. (9). For example, consider a continuous-time deterministic EnKBF formulation (Sakov and Oke, 2008; Bergemann and Reich, 2012):
$\mathrm{d}\vec{x}^{(i)}_{\mathrm{f}}=\mathbf{F}\vec{x}^{(i)}_{\mathrm{f}}\mathrm{d}t+\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)}+\mathbf{P}_{\mathrm{f}}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\left(\mathrm{d}\vec{y}-\frac{\mathbf{H}\vec{x}^{(i)}_{\mathrm{f}}+\mathbf{H}\overline{\vec{x}}_{\mathrm{f}}}{2}\mathrm{d}t\right).$ (40)
Its anomalies obey the dynamics with a half-gain drift (cf. Eq. (35)):
$\mathrm{d}\vec{\hat{x}}_{\mathrm{f}}^{(i)}=\left(\mathbf{F}-\frac{1}{2}\mathbf{K}\mathbf{H}\right)\vec{\hat{x}}_{\mathrm{f}}^{(i)}\mathrm{d}t+\mathbf{\Sigma}^{1/2}\mathrm{d}\vec{B}^{(i)}.$ (41)
@@
Clearly, Eq. (42) does not match the continuous-time Riccati equation Eq. (37) for the filter covariance $\mathbf{P}_{\mathrm{f}}$. To see the discrepancy explicitly, we define the difference matrix $\mathbf{D}:=\mathbf{P}_{\mathrm{f},\mathrm{s}}-\mathbf{P}_{\mathrm{f}}$. Subtracting Eq. (37) from Eq. (42) yields
@@
Although $\mathbf{D}(T)=\mathbf{0}$, the positive semi-definite term $\frac{1}{2}\mathbf{P}_{\mathrm{f}}\mathbf{H}^{\top}\mathbf{\Gamma}^{-1}\mathbf{H}\mathbf{P}_{\mathrm{f}}$ generally forces $\mathbf{D}(t)\neq\mathbf{0}$ for $t<T$. Hence, $\mathbf{P}_{\mathrm{f},\mathrm{s}}\neq\mathbf{P}_{\mathrm{f}}$ for this deterministic EnKBF. Consequently, the resulting smoother ensemble covariance $\mathbf{P}_{\mathrm{s}}$ will deviate from the RTS smoother covariance equation Eq. (39).
Nevertheless, despite this theoretical mismatch, we have tested this deterministic EnKBF Eq. (40) as a forward pass in our framework. It yields results similar to those of the stochastic EnKBF Eq. (7) across all numerical experiments presented in this paper (not shown). In addition, for either choice of the forward EnKBF, the smoother consistently improves upon the filter by reducing both the estimation error and the ensemble spread (with behavior similar to that shown in Figure 6).
[…]
NC and MA conceived the project, with NC designing the numerical experiments. ZJ developed the code, performed experiments, and prepared the manuscript draft and figures, in regular discussion with NC and MA. SR reviewed an earlier version and provided critical feedback that motivated the analysis in Appendix B. NC supervised the project, and all authors contributed to manuscript revision.
[…]
NC is a member of the editorial board of Nonlinear Processes in Geophysics. The authors have no other competing interests to declare.
Acknowledgements.
The research of NC is funded by the Army Research Office W911NF-23-1-0118. ZJ and MA are partially supported as research assistants under this grant. MA is also supported in part by NSF Award DMS-2023239. The work of SR has been partially funded by Deutsche Forschungsgemeinschaft (DFG) – Project-ID 318763901 – SFB1294.